OpenStudy (anonymous):

5+2i/-6-2i in the form a+bi

OpenStudy (triciaal):

rationalize the denominator this means multiply by 1 using the denominator and switch the sign of the 2nd term for this problem denominator is (-6-2i) so we multiply by (-6+2i) 

OpenStudy (anonymous):

Thanks, but I'm still not able to come up with the answer that matches the key which is -17-i/20??

OpenStudy (triciaal):

can you show your work? i'll also do it

OpenStudy (anonymous):



OpenStudy (triciaal):

you just have calculation error somewhere the answer is correct the form is -17/20-i/20 remember to watch your signs I^2 = -1 neg times neg is positive

OpenStudy (anonymous):

I appreciate your help - I just don't get it. Thanks for trying to help me.

OpenStudy (triciaal):

sorry I have computer problems and could not repond for the numerator 2i*2i = 4 I^2 and I^2= -1 so it becomes -4 the numerator is therefore -30-4 -2i = -34-2i the denominator 36-4i^2 = 36 + 4 = 40 reduce to the lowest term by dividing by 2 and write in the form a + bi (-34-2i) /40 = -17/20 -I/20